# Recurrent Self-Attention Dynamics: An Energy-Agnostic Perspective from Jacobians

### v2

@@ 5.3 Lyapunov exponent indicates criticality
The Lyapunov exponent measures the exponential rate at which trajectories locally converge or diverge in dynamical systems (Khalil, 2002) (see details in Section B.3). Intuitively, it corresponds to the time-averaged value of $\ln\sigma_{i}(\bm{J})$, where $\sigma_{i}(\bm{J})$ denotes the Jacobian’s singular values. A positive exponent indicates instability and sensitivity to initial conditions, whereas a negative exponent implies convergence. An exponent close to zero characterizes a critical regime, often referred to as the edge of chaos, where signals neither explode nor vanish and can propagate for a long period. This critical regime has been reported to correlate with the high performance of neural networks across various contexts (Boedecker et al., 2012; Poole et al., 2016; Pennington et al., 2017).
Figure 3 shows that SA models achieving high test accuracy empirically have Lyapunov exponents close to zero, thus operating near criticality. The maximum and mean exponents display nearly identical behaviors. We vary hyper-parameters, including $\eta$ and the norm of weight matrices, and plot multiple points (see details in Section B.3). In models with normalization layers, the exponent tends to concentrate around zero, indicating the criticality. High accuracy is achieved only in these models. This supports the stabilizing effect of normalization implied by Proposition 5.1 and Section 5.2. In contrast, energy-based symmetric SA models show negative exponents, consistent with their Lyapunov stability, although these constraints reduce their trainability, leading to lower accuracy. Interestingly, the maximum Lyapunov exponent of successful models is slightly positive ($\sim 0.1$), indicating that the dynamical state resides near criticality from the chaotic side. This observation aligns with previous reports of positive maximum Lyapunov exponents in some Transformers  (Inoue et al., 2022; Liu et al., 2024; Tong et al., 2025). We also confirmed similar Lyapunov exponents on the CIFAR-10 dataset (Section B.6).
@@ 6 Quantitative insight into inference dynamics
Here, we experimentally investigate the test-time scaling of inference in looped architectures. We mainly focus on evaluation on the Sudoku task using the SATNet (Wang et al., 2019) dataset for in-distribution (ID) data and the RRN dataset (Palm et al., 2018) for out-of-distribution (OOD) data. At test time, we increased the number of loops beyond the training setting of $T=16$. Details are provided in Appendix B.
@@ 6.1 Test-time scaling and normalization
Miyato et al. (2025) showed that AKOrN exhibits test-time scaling, whereas ItrSA does not, suggesting the superiority of AKOrN over ItrSA. However, with our formulation of ItrSA, Figure 4 demonstrates that ItrSA also exhibits test-time scaling. Moreover, it shows that AKOrN fails to maintain test-time scaling when $N$ becomes large, which is consistent with the observations by Miyato et al. (2025). This issue can be mitigated by applying RMS normalization, where we use oscillator-wise RMS normalization. The learned scaling parameter $\bm{\gamma}$ prevents the Jacobian from becoming excessively large. Empirically, we confirmed that the trained $\bm{\gamma}$ remains small (see Table S.2) eliminating the need for explicit clipping such as $|\gamma_{i}|\leq 1$.
Figure 4: ItrSA consistently improves accuracy as the number of loops $T$ increases, regardless of the value of oscillator dimension $N$. Note that $N=512$ corresponds to the setting where tokens are not split into oscillators. Error bars indicate the standard deviation.
@@ 6.2 Application to regularization
Method (i): Energy-based regularization. Proposition 4.2 suggests that SA architectures satisfying specific conditions inherently minimize an energy function. To investigate the practical utility of this property, we consider applying the constraints identified in the proposition as a regularization term to the existing multi-head MSA. We add the following energy-based regularization term to the loss function during the training of ItrSA, which can be interpreted as an approximation of energy-based SA models. By defining the concatination as $\bm{W}^{V}\coloneqq[\bm{W}_{1}^{V},\cdots,\bm{W}_{H}^{V}]\in\mathbb{R}^{D\times D}$, we introduce
@@
where both $\bm{W}^{V}$ and $\bm{W}^{O}$ are implemented as orthogonal matrices. Note that each $\bm{W}_{h}^{V}$ in the proposition is interpreted as the product $\bm{W}_{h}^{V}\bm{W}_{h}^{O}$ in ItrSA. If $R_{\text{E-multi}}=0$, $\bm{W}_{h}^{V}\bm{W}_{h}^{O}$ becomes symmetric, as implied by Proposition 4.2. For a single-head case, we can also propose Proposition 4.1 as a regularization term $R_{\text{E-single}}$ (see Section B.2).
@@
Figure 5 shows the effects of the regularization methods. The accuracy of E-multi is lower than that without regularization. These results suggest that energy-based regularization may be unnecessary, casting doubt on the validity of the energy-based perspective for practical applications. E-single fails to reduce the training loss and fails even on ID tasks possibly due to the single-head constraint.
@@
Spectral regularization substantially enhances the performance of both AKOrN and ItrSA, and the effect is especially pronounced in AKOrN. Eq. (15) shows that the maximum eigenvalue of the Jacobian can grow significantly when the weight matrices have large norms. Spectral regularization addresses this issue by directly constraining the Jacobian spectrum.
Figure 5: Energy-based regularization (“E-single” and “E-multi”) underperforms the original methods, while Jacobian spectral regularization (“Spec”) overperforms. Error bars indicate standard deviation.
@@ 6.3 An interpretation of pseudo-energy via Jacobian
Figure 6: Test-time inference of AKOrN on the Sudoku dataset.
We found that the Jacobian provides an interpretation. Suppose that $\operatorname{vec}(\operatorname{MSA}(\bm{X}^{(t)}))\in\mathbb{R}^{DS}$ is well-approximated by $\bm{J}_{t}\bm{x}_{t}$, where $\bm{x}_{t}\in\mathbb{R}^{DS}$ denotes $\operatorname{vec}(\bm{X}^{(t)})$ and $\bm{J}_{t}=\partial\operatorname{MSA}(\bm{X}^{(t)})/\partial\bm{X}^{(t)}\in\mathbb{R}^{DS\times DS}$. Figure 6 shows that while $E_{\text{pseudo}}$ significantly decreases, the cosine similarity between $\operatorname{vec}(\operatorname{MSA}(\bm{X}^{(t)}))$ and $\bm{J}_{t}\bm{x}_{t}$ remains high throughout iterations. Under this approximation, neglecting a constant term $\bm{C}$ in $\Delta\bm{X}$, we obtain the relation $E_{\text{pseudo}}=-\bm{x}_{t}^{\top}\bm{S}_{t}\bm{x}_{t}/2$, a quadratic form involving a symmetric matrix $\bm{S}_{t}=\bm{J}_{t}+\bm{J}_{t}^{\top}$. We further expand $\bm{x}_{t}$ in the orthonormal basis of $\bm{S}_{t}$ as $\bm{x}_{t}=\sum_{k=1}^{DS}a_{t}^{k}\bm{v}_{t}^{k}$, and the eigenvectors are sorted in descending order of their eigenvalues ($\lambda_{t}^{1}\geq\dots\geq\lambda_{t}^{DS}$). The contribution index in the figure quantifies the extent to which the state $\bm{x}_{t}$ is captured by the top $2\%$ eigenvalues, specifically $\sum_{k\leq 0.02DS}(a_{t}^{k})^{2}/\sum_{k=1}^{DS}(a_{t}^{k})^{2}$. During inference, the proportion of the state in the top $2\%$ eigenspace increases monotonically, eventually dominating nearly $40\%$. This suggests that the loop effectively performs a computation analogous to power iteration, but constrained to real space with positive eigenvalues. As a side note, in Section B.5, a more detailed analysis of the Jacobian matrix implies that this alignment to the eigenspace is predominantly determined by a certain time-independent matrix. Thus, the Jacobian provides a meaningful and promising interpretation for the observed decrease of pseudo-energy during AKOrN inference.
@@ 7 Conclusion
In this paper, we advanced the understanding and control of recurrent self-attention (SA) from a dynamical systems perspective. First, we generalized energy-based formulations to weaker symmetry and multi-head configurations, closer to realistic settings. However, experiments showed that energy-regularized SA underperformed standard SA, suggesting other factors at play in practical models. Second, we analyzed the Jacobian matrix of standard SA architectures, revealing that normalization layers effectively regularize their spectral properties. We further clarify that Lyapunov exponents at criticality characterize high-performance inference, which indicates that the state update of SA is more dynamic than energy-constrained cases. We also argued how Jacobians provide quantitative insights into performance-enhancing regularization and pseudo-energy behaviors.
@@
In this work, we focused on recurrent SAs without positional encoding, masking, or MLP blocks, unlike looped Transformers in practice. Investigating how these components alter state dynamics remains an interesting direction. Regarding theoretical limitations, the upper bound provided by Proposition 4.1 provides a conservative upper bound that is looser than empirical observations, an issue also noted in existing analyses of SA’s Jacobians without normalization. Deriving tighter bounds remains an intriguing open theoretical problem. Additionally, analytically justifying empirical phenomena, such as the Lyapunov exponent concentration (Section 5.3) and the Jacobian-based approximation (Section 6.1), by solving state dynamics would be challenging yet exciting themes for theory. We expect our findings to serve as a foundation for further theoretical and practical advancements in looped SA architectures and their rich inference dynamics.
Appendices
Appendix A Proof
@@ A.1 Lemmas
We use the following lemma in our derivation.
@@
the result follows from lemma A.2. ∎
@@
Proposition 4.2 is restated. Consider the continuous-time dynamics for multi-head SA without projection: $d\bm{X}/dt=\sum_{h=1}^{H}SA_{h}(\bm{X})$. An energy function
@@
Proposition 5.1 is restated. Suppose that, in the update of ItrSA (9), the input to the normalization layer satisfies $\|\bm{X}_{[i,:]}+\eta\Delta\bm{X}_{[i,:]}\|\geq R$ for all $i\in[1,S]$. Then, the spectral norm of the Jacobian satisfies the upper bound
$\left\|\frac{\partial\operatorname{RMSNorm}(\bm{X}+\eta\Delta\bm{X})}{\partial\bm{X}}\right\|_{2}$ $\leq\frac{\max_{j}(|\gamma_{j}|)}{R}\left(1+|\eta|\left\|\bm{J}_{\text{MSA}}(\bm{X})\right\|_{2}\right),$ (67)
@@
First, for any vector $a\in\mathbb{R}^{D}$, the eigenvalues of the matrix $I_{D}-\frac{aa^{\top}}{\|a\|^{2}}$ are $1$ (with multiplicity $D-1$) and 00 (with multiplicity $1$). Hence,
$\left\|I_{D}-\frac{aa^{\top}}{\|a\|^{2}}\right\|_{2}$ $\leq 1.$ (68)
Using Lemma A.3, we have
@@
$\leq\frac{\max_{j}(|\gamma_{j}|)}{R}\left(1+|\eta|\left\|\bm{J}_{\text{MSA}}(\bm{X})\right\|_{2}\right).$ (76)
@@
where we used the fact that for an antisymmetric matrix $\bm{\Omega}$, $\bm{x}^{\top}\bm{\Omega}\bm{x}=0$. Note also that an antisymmetric matrix has eigenvalues of the form ±$\pm i\omega_{j}$, where $\omega_{j}\geq 0$ ($j=1,2,\dots$). For simplicity, assume all eigenvalues have identical magnitude $\omega_{j}=\omega$. Then, we have

### v3

@@ 5.3 Lyapunov exponent indicates criticality
The Lyapunov exponent measures the exponential rate at which trajectories locally converge or diverge in dynamical systems (Khalil, 2002) (see details in Section˜B.3). Intuitively, it corresponds to the time-averaged value of $\ln\sigma_{i}(\bm{J})$, where $\sigma_{i}(\bm{J})$ denotes the Jacobian’s singular values. A positive exponent indicates instability and sensitivity to initial conditions, whereas a negative exponent implies convergence. An exponent close to zero characterizes a critical regime, often referred to as the edge of chaos, where signals neither explode nor vanish and can propagate for a long period. This critical regime has been reported to correlate with the high performance of neural networks across various contexts (Boedecker et al., 2012; Poole et al., 2016; Pennington et al., 2017).
Figure˜3 shows that SA models achieving high test accuracy empirically have Lyapunov exponents close to zero, thus operating near criticality. The maximum and mean exponents display nearly identical behaviors. We vary hyper-parameters, including $\eta$ and the norm of weight matrices, and plot multiple points (see details in Section˜B.3). In models with normalization layers, the exponent tends to concentrate around zero, indicating the criticality. High accuracy is achieved only in these models. This supports the stabilizing effect of normalization implied by Proposition˜5.1 and Section 5.2. In contrast, energy-based symmetric SA models show negative exponents, consistent with their Lyapunov stability, leading to lower accuracy. Interestingly, the maximum Lyapunov exponent of successful models is slightly positive ($\sim 0.1$), indicating that the dynamical state resides near criticality from the chaotic side. Notably, we observed that the dynamics with this slightly positive maximum Lyapunov exponent indicate the sensitivity to initial conditions, implying chaotic behavior (see Figure˜S.5 in the appendix). This observation aligns with previous reports of positive maximum Lyapunov exponents in some Transformers  (Inoue et al., 2022; Liu et al., 2024; Tong et al., 2025). We observed similar Lyapunov exponents across the CIFAR-10 dataset (Figure˜S.7) and the language modeling task (Table˜S.4). We further observed that as the number of attention heads increases, the Lyapunov exponents tend to increase (Figures˜S.2 and S.1). This implies that multi-head attention would favor a more dynamic state.
@@ 6 Quantitative insight into inference dynamics
Here, we experimentally investigate the test-time scaling of inference in looped architectures. We mainly focus on evaluation on the Sudoku task using the SATNet (Wang et al., 2019) dataset for in-distribution (ID) data and the RRN dataset (Palm et al., 2018) for out-of-distribution (OOD) data. At test time, we increased the number of loops beyond the training setting of $T=16$. Details are provided in Appendix˜B.
@@ 6.1 Test-time scaling and normalization
Miyato et al. (2025) showed that AKOrN exhibits test-time scaling, whereas ItrSA does not, suggesting the superiority of AKOrN over ItrSA. However, with our formulation of ItrSA, Figure˜4 demonstrates that ItrSA also exhibits test-time scaling. Moreover, it shows that AKOrN fails to maintain test-time scaling when $N$ becomes large, which is consistent with the observations by Miyato et al. (2025). This issue can be mitigated by applying RMS normalization, where we use oscillator-wise RMS normalization. The learned scaling parameter $\bm{\gamma}$ prevents the Jacobian from becoming excessively large. Empirically, we confirmed that the trained $\bm{\gamma}$ remains small (see Table˜S.3), eliminating the need for explicit clipping such as $|\gamma_{i}|\leq 1$.
Figure 4: ItrSA consistently improves accuracy as the number of loops $T$ increases, regardless of the value of oscillator dimension NN. Note that $N=512$ corresponds to the setting where tokens are not split into oscillators. Error bars indicate the standard deviation.
@@ 6.2 Application to regularization
To apply our energy-based and Jacobian-based analysis, we add the following regularization term $R$, scaled by a tunable parameter $\lambda$.
Method (i): Energy-based regularization. Proposition˜4.2 suggests that SA architectures satisfying specific conditions inherently minimize an energy function. To investigate the practical utility of this property, we consider applying the constraints identified in the proposition as a regularization term to the existing multi-head MSA. We add the following energy-based regularization term to the loss function during the training of ItrSA, which can be interpreted as an approximation of energy-based SA models. By defining the concatenation as $\bm{W}^{V}\coloneqq[\bm{W}_{1}^{V},\cdots,\bm{W}_{H}^{V}]\in\mathbb{R}^{D\times D}$, we introduce
@@
where both $\bm{W}^{V}$ and $\bm{W}^{O}$ are implemented as orthogonal matrices. Note that each $\bm{W}_{h}^{V}$ in the proposition is interpreted as the product $\bm{W}_{h}^{V}\bm{W}_{h}^{O}$ in ItrSA. If $R_{\text{E-multi}}=0$, $\bm{W}_{h}^{V}\bm{W}_{h}^{O}$ becomes symmetric as implied by Proposition˜4.2. For a single-head case, we can also propose Proposition˜4.1 as a regularization term $R_{\text{E-single}}$ (see Section˜B.2).
@@
Figure˜5 shows the effects of the regularization methods. The accuracy of E-multi is lower than that without regularization. The Lyapunov exponents shown in  Figure˜S.2 indicates that multi-head energy regularization encourages more convergent dynamics and this does not necessarily yield better performance. E-single fails to reduce the training loss and fails even on ID tasks possibly due to the single-head constraint. These results suggest that energy-based regularization may be unnecessary, casting doubt on the validity of the energy-based perspective for practical applications.
@@
Spectral regularization substantially enhances the performance of both AKOrN and ItrSA, and the effect is especially pronounced in AKOrN. Equation (16) shows that the maximum eigenvalue of the Jacobian can grow significantly when the weight matrices have large norms. Spectral regularization addresses this issue by directly constraining the Jacobian spectrum. When $N=8$, AKOrN achieves the best performance and spectral regularization is also effective (Figure˜S.3).
Figure 5: Energy-based regularization (“E-single” and “E-multi”) underperforms the original methods, while Jacobian spectral regularization (“Spec”) outperforms. We set $H=8$ except for E-single ($H=1$). For AKOrN, we used $N=4$.
@@ 6.3 An interpretation of pseudo-energy via Jacobian
Figure 6: Test-time inference of AKOrN on the Sudoku dataset.
@@
We found that the Jacobian provides an interpretation. Suppose that $\operatorname{vec}(\operatorname{MSA}(\bm{X}^{(t)}))\in\mathbb{R}^{DS}$ is well-approximated by $\bm{J}_{t}\bm{x}_{t}$, where $\bm{x}_{t}\in\mathbb{R}^{DS}$ denotes $\operatorname{vec}(\bm{X}^{(t)})$ and $\bm{J}_{t}=\partial\operatorname{MSA}(\bm{X}^{(t)})/\partial\bm{X}^{(t)}\in\mathbb{R}^{DS\times DS}$. Figure˜6 shows that while $E_{\text{pseudo}}$ significantly decreases, the cosine similarity between $\operatorname{vec}(\operatorname{MSA}(\bm{X}^{(t)}))$ and $\bm{J}_{t}\bm{x}_{t}$ remains high throughout iterations. Under this approximation, neglecting a constant term $\bm{C}$ in $\Delta\bm{X}$, we obtain the relation $E_{\text{pseudo}}=-\bm{x}_{t}^{\top}\bm{S}_{t}\bm{x}_{t}/2$, a quadratic form involving a symmetric matrix $\bm{S}_{t}=\bm{J}_{t}+\bm{J}_{t}^{\top}$. We further expand $\bm{x}_{t}$ in the orthonormal basis of $\bm{S}_{t}$ as $\bm{x}_{t}=\sum_{k=1}^{DS}a_{t}^{k}\bm{v}_{t}^{k}$, and the eigenvectors are sorted in descending order of their eigenvalues ($\lambda_{t}^{1}\geq\dots\geq\lambda_{t}^{DS}$). The contribution index in the figure quantifies the extent to which the state $\bm{x}_{t}$ is captured by the top $2\%$ eigenvalues, specifically $\sum_{k\leq 0.02DS}(a_{t}^{k})^{2}/\sum_{k=1}^{DS}(a_{t}^{k})^{2}$. During inference, the proportion of the state in the top $2\%$ eigenspace increases monotonically, eventually dominating nearly $40\%$. This suggests that the loop effectively performs a computation analogous to power iteration, but constrained to real space with positive eigenvalues. Thus, the Jacobian provides a meaningful and promising interpretation for the observed decrease of pseudo-energy during AKOrN inference.
As a side note, in Section B.5, a more detailed analysis of the Jacobian matrix implies that this alignment to the eigenspace is predominantly determined by a certain time-independent matrix.
@@ 7 Conclusion
In this paper, we advanced the understanding and control of recurrent self-attention (SA) from a dynamical systems perspective. First, we generalized energy-based formulations to weaker symmetry and multi-head configurations closer to realistic settings. However, experiments showed that energy-regularized SA underperformed standard SA, suggesting other factors at play in practical models. Second, we analyzed the Jacobian matrix of standard SA architectures, revealing that normalization layers effectively regularize their spectral properties. We further clarify that Lyapunov exponents at criticality characterize high-performance inference, which indicates that the state update of SA is more dynamic than energy-constrained cases. We also argued how Jacobians provide quantitative insights into performance-enhancing regularization and pseudo-energy behaviors.
@@
In this work, we focused on recurrent SAs without positional encoding, masking, or MLP blocks, unlike looped Transformers in practice. Investigating how these components alter state dynamics remains an interesting direction. Regarding theoretical limitations, the upper bound in Proposition˜5.1 provides a conservative upper bound that is looser than empirical observations, an issue also noted in existing analyses of SA’s Jacobians without normalization. Deriving tighter bounds remains an intriguing open theoretical problem. Additionally, analytically justifying empirical phenomena, such as the Lyapunov exponent concentration (Section˜5.3) and the Jacobian-based approximation (Section˜6.3), by solving state dynamics would be challenging yet exciting themes for theory. We expect our findings to serve as a foundation for further theoretical and practical advancements in looped SA architectures and their rich inference dynamics.
Acknowledgments and Disclosure of Funding
We thank the reviewers for their insightful and helpful feedback on the manuscript, and Takeru Miyato for his valuable comments. The authors acknowledge the funding support from JST FOREST (Grant No. JPMJFR226Q). RK is also supported by JSPS KAKENHI (Grant Nos. 22H05116, 23K16965).
Appendix A Proofs
@@ A.1 Lemmas
We use the following lemmas in our derivation.
@@
the result follows from lemma˜A.2. ∎
@@
Proposition˜4.2 is restated. Consider the continuous-time dynamics for multi-head SA without projection: $d\bm{X}/dt=\sum_{h=1}^{H}SA_{h}(\bm{X})$. An energy function
@@
Proposition˜5.1 is restated. Suppose that, in the update of ItrSA (9), the input to the normalization layer satisfies $\|\bm{X}_{[i,:]}+\eta\Delta\bm{X}_{[i,:]}\|\geq R$ for all $i\in[1,S]$. Then, the spectral norm of the Jacobian satisfies the upper bound
$\left\|\frac{\partial\operatorname{RMSNorm}(\bm{X}+\eta\Delta\bm{X})}{\partial\bm{X}}\right\|_{2}$ $\leq\frac{\max_{j}(|\gamma_{j}|)}{R}\left(1+\eta\left\|\bm{J}_{\text{MSA}}(\bm{X})\right\|_{2}\right),$ (68)
@@
First, for any vector $a\in\mathbb{R}^{D}$, the eigenvalues of the matrix $I_{D}-\frac{aa^{\top}}{\|a\|^{2}}$ are $1$ (with multiplicity D1D-1) and 0 (with multiplicity $1$). Hence,
$\left\|I_{D}-\frac{aa^{\top}}{\|a\|^{2}}\right\|_{2}$ $=1.$ (69)
Using Lemma˜A.3, we have
@@
$\leq\frac{\max_{j}(|\gamma_{j}|)}{R}\left(1+\eta\left\|\bm{J}_{\text{MSA}}(\bm{X})\right\|_{2}\right).$ (77)
Here, we can also show that, in the limit $\eta\to\infty$, the Jacobian norm remains $O(1)$. Define, for each $i$, the projection
$\bm{P}_{i}\coloneqq\bm{I}_{D}-\frac{\bm{Y}_{[i,:]}\bm{Y}_{[i,:]}^{\top}}{\|\bm{Y}_{[i,:]}\|_{2}^{2}},$ (78)
so that $\|\bm{P}_{i}\|_{2}=1$. Let $\bm{D}\coloneqq\operatorname{diag}(\bm{\gamma})$, and define the block factors
$\bm{A}_{i}(\eta)\coloneqq\frac{1}{\|\bm{X}_{[i,:]}+\eta\,\Delta\bm{X}_{[i,:]}\|_{2}}\,\bm{D}\,\bm{P}_{i},\qquad\bm{A}(\eta)\coloneqq\operatorname{blockdiag}\left(\{\bm{A}_{i}(\eta)\}_{i=1}^{N}\right).$ (79)
By the triangle inequality, for each $i$,
$\|\bm{X}_{[i,:]}+\eta\,\Delta\bm{X}_{[i,:]}\|_{2}\;\geq\;\eta\,\|\Delta\bm{X}_{[i,:]}\|_{2}-\|\bm{X}_{[i,:]}\|_{2}.$ (80)
Recall that $\bm{X}$ is the output of the previous layer, so that
$\bm{X}=\operatorname{RMSNorm}(\bm{Z})$ (81)
for some $\bm{Z}\in\mathbb{R}^{S\times D}$. Therefore,
$\|\bm{X}_{[i,:]}\|_{2}$ $=\|\operatorname{RMSNorm}(\bm{Z})_{[i,:]}\|_{2}$ (82)
$=\Big\|\operatorname{diag}(\bm{\gamma})\,\frac{\bm{Z}_{[i,:]}}{\|\bm{Z}_{[i,:]}\|_{2}}\Big\|_{2}$ (83)
$=\frac{\sqrt{\sum_{j=1}^{D}\gamma_{j}^{2}Z_{[i,j]}^{2}}}{\|\bm{Z}_{[i,:]}\|_{2}}$ (84)
$\leq\max_{j}|\gamma_{j}|\,\frac{\sqrt{\sum_{j=1}^{D}Z_{[i,j]}^{2}}}{\|\bm{Z}_{[i,:]}\|_{2}}$ (85)
$=\max_{j}|\gamma_{j}|$ (86)
for each $i=1,\ldots,S$.
Assume that $\min_{i}\|\Delta\bm{X}_{[i,:]}\|_{2}\geq\varepsilon$ for some constant $\varepsilon>0$. Then, for a sufficiently large $\eta$ satisfying $\eta\geq\max_{j}|\gamma_{j}|/\varepsilon$, we have
$\eta\geq 2\,\frac{\|\bm{X}_{[i,:]}\|_{2}}{\|\Delta\bm{X}_{[i,:]}\|_{2}}.$ (87)
Hence,
$\|\bm{X}_{[i,:]}+\eta\,\Delta\bm{X}_{[i,:]}\|_{2}$ $\geq\eta\,\|\Delta\bm{X}_{[i,:]}\|_{2}-\|\bm{X}_{[i,:]}\|_{2}$ (88)
$\geq\tfrac{\eta}{2}\,\|\Delta\bm{X}_{[i,:]}\|_{2}.$ (89)
Using submultiplicativity and $\|\bm{P}_{i}\|_{2}=1$, we obtain
$\|\bm{A}_{i}(\eta)\|_{2}$ $=\frac{\|\bm{D}\,\bm{P}_{i}\|_{2}}{\|\bm{X}_{[i,:]}+\eta\,\Delta\bm{X}_{[i,:]}\|_{2}}$ (90)
$\leq\frac{\|\bm{D}\|_{2}}{\|\bm{X}_{[i,:]}+\eta\,\Delta\bm{X}_{[i,:]}\|_{2}}$ (91)
$\leq\frac{2\,\|\bm{D}\|_{2}}{\eta\,\|\Delta\bm{X}_{[i,:]}\|_{2}},$ (92)
for all $i=1,\ldots,S$.
For a block-diagonal matrix, the operator norm equals the maximum block norm; thus
$\|\bm{A}(\eta)\|_{2}$ $=\max_{i}\|\bm{A}_{i}(\eta)\|_{2}$ (93)
$\leq\max_{i}\frac{2\,\|\bm{D}\|_{2}}{\eta\,\|\Delta\bm{X}_{[i,:]}\|_{2}}$ (94)
$=\frac{2\,\|\bm{D}\|_{2}}{\eta\,\min_{i}\|\Delta\bm{X}_{[i,:]}\|_{2}}$ (95)
$\leq\frac{2\,\|\bm{D}\|_{2}}{\eta\,\varepsilon}.$ (96)
Therefore, we have
$\left\|\bm{A}(\eta)\,(\bm{I}_{ND}+\eta\,\bm{J}_{\mathrm{MSA}})\right\|_{2}$ $\leq\|\bm{A}(\eta)\|_{2}\left(\|\bm{I}_{ND}\|_{2}+\eta\,\|\bm{J}_{\mathrm{MSA}}\|_{2}\right)$ (97)
$\leq\frac{2\,\|\bm{D}\|_{2}}{\varepsilon}\left(\|\bm{J}_{\mathrm{MSA}}\|_{2}+\frac{1}{\eta}\right).$ (98)
Since Eq. (16) makes a constant upper-bound of $\|\bm{J}_{\mathrm{MSA}}\|_{2}$, for $\eta\to\infty$, we obtain
$\left\|\frac{\partial\,\operatorname{RMSNorm}(\bm{X}+\eta\,\Delta\bm{X})}{\partial\bm{X}}\right\|_{2}=O(1).$ (99)
@@
where we used the fact that for an antisymmetric matrix $\bm{\Omega}$, $\bm{x}^{\top}\bm{\Omega}\bm{x}=0$. Note also that an antisymmetric matrix has eigenvalues of the form $\pm i\omega_{j}$, where $\omega_{j}\geq 0$ ($j=1,2,\dots$). For simplicity, assume all eigenvalues have identical magnitude $\omega_{j}=\omega$. Then, we have